# AnoVox: A Benchmark for Multimodal Anomaly Detection in Autonomous Driving

### v3

@@ Abstract
The scale-up of autonomous vehicles depends heavily on their ability to deal with anomalies, such as rare objects on the road. In order to handle such situations, it is necessary to detect anomalies in the first place. Anomaly detection for autonomous driving has made great progress in the past years but suffers from poorly designed benchmarks with a strong focus on camera data. In this work, we propose AnoVox, the largest benchmark for ANOmaly detection in autonomous driving to date. AnoVox incorporates large-scale multimodal sensor data and spatial VOXel ground truth, allowing for the comparison of methods independent of their used sensor. We propose a formal definition of normality and provide a compliant training dataset. AnoVox is the first benchmark to contain both content and temporal anomalies.
@@ 1 Introduction
In the past years, we have seen autonomous vehicles scale from small and well-defined Operational Design Domains (ODD) to city-wide areas, potentially reaching geographic areas covering multiple neighboring cities in the near future [73, 75, 74]. However, with a growing fleet in an expanding ODD, scenarios from the long tail of rare events [71] occur more frequently as a consequence of increased exposure. Expert perspectives, focusing on content and temporal anomalies, are commonly used to judge "long tail" data points [18, 17, 40, 10, 11, 61, 66], as replaying training data during inference is not possible. Such expert views are often used for the creation of anomaly detection benchmarks but are of no help for anomaly detection. Few publications take a data-oriented perspective on the definition of anomalies [39, 25, 82], also known as out-of-distribution (OOD), corner cases, or outliers.
Recent research has focused on anomaly detection methods [42] and its ramifications, such as OOD segmentation [58], object detection [31], instance segmentation [55], and video tracking [54]. These methods typically do not learn a representation of normality from training data alone, but require a semantic segmentation stage and fine-tuning on auxiliary data containing (artificial) anomalies [9, 23]. While there exists a considerable number of datasets in autonomous driving [53, 49, 14], only around 5 - 10% of datasets are designed for benchmarking anomaly detection methods in autonomous driving [15, 13]. These existing datasets, however, have significant limitations. First, the majority are camera-only, despite autonomous driving relying on multi-modal sensor setups. Second, temporal information is often overlooked, limiting anomaly detection to single frames. Additionally, content anomalies are rarely found in challenging traffic situations, which allows for the assumption that anything on the road might be an anomaly [54]. Most anomalies are human-defined, such as dogs on the street [9], making detection challenging as these classes are typically included in training datasets but not annotated [27, 31]. As a result, anomaly detection methods can miss anomalies they don’t perceive as atypical due to their training.
@@
With AnoVox, we offer a challenging and adaptable benchmark for ANOmaly detection, leveraging the CARLA simulation engine [30] to provide ground truth data in all modalities and spatial VOXel representation. Given the surge in foundation models [21, 60], world models [80, 45], and large-scale pre-training approaches [43, 78], we see potential for the application of self- and unsupervised training paradigms in the field of anomaly detection as well. Hence, we prioritize support for label-free training paradigms and provide an extensive synthetic dataset to enable research in OOD segmentation, addressing the limitations of current datasets and enabling rapid prototyping for technologies that fill the blind spots of current anomaly detection methods. Our contributions include:
@@
All code to reproduce our results, create new AnoVox instantiations, and benchmark anomaly detection methods can be found on GitHub. The dataset can be found on Zenodo.
@@ 2 Related Work
In autonomous driving, anomaly detection has been classically evaluated on a small set of benchmarks. For our analysis of related works, we have included works with published open-access perception datasets from an ego perspective that provide pixel- or point-wise ground truth. Thus, we neglected frameworks that do not provide explicit data [77, 68, 38, 76, 65], such with missing or incomplete data [34, 52, 48], and any that provide only frame-wise annotations [67, 79, 70, 2]. For our overview of related works as shown in Table 1, we updated and extended the survey of Bogdoll et al. [15], comparing now 22 datasets from the literature.
@@
There are different categories of how anomalies were introduced [15] in the datasets. Recording and Simulation are similar in the way that selected anomalies were directly introduced into the data. This way, the anomalies are truly part of the environment [62, 19, 54]. The definition of what counts as an anomaly can vary, though. Data Augmentation typically follows a copy-and-paste pattern, where images of anomalies are pasted onto an already existing scene from another dataset. This way, a distribution shift between the anomaly and the underlying data is introduced [9, 6]. Web Sourcing describes the process of manually curating images from the web that are deemed anomalous [23, 46]. Class Exclusion is based on existing datasets and removes selected classes from the training data, thus treating them as anomalous, while the classes themselves remain rather normal from a human point of view [35, 41]. Void Classes utilizes void or misc classes from existing datasets and labels them as anomalous. This can be done with additional labeling guidelines to only relabel selected ones [51]. Finally, OOD Detection uses an anomaly detection method to derive anomaly proposals from a dataset which can then be labeled, typically after a human quality inspection [51].
@@
AnoVox is a challenging benchmark that addresses the field of anomaly detection from an embodied AI perspective rather than from a pure computer vision perspective. Therefore, we provide data in a temporal context with state information about the ego vehicle for typical multimodal sensor setups. For a clear definition of what anomalies are, we provide a formalized definition of normality and provide compliant training data. This especially allows for the usage of un- and self-supervised anomaly detection methods, which often rely on large amounts of unlabeled data.
@@ 3.2 Scenario Generation
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 4: Content anomalies:. Examples from our six categories: An ape as an animal anomaly, an old tv as a home anomaly, a statue as a special anomaly, a tree as a nature anomaly, a pillar as a falling anomaly, and a hot air balloon as an airplane anomaly.
@@
We provide 178 different content anomalies in five different size classes tiny, small, medium, big, and huge. Semantically, they hierarchically belong to six different super-classes. Every anomaly has an individual label next to its super-class and size label.
The class animal includes 33 animals of different sizes. The category home includes 53 typical household items such as furniture, tables, backpacks, or cardboard boxes. The category special includes 67 objects of rather atypical types and such that fall into a misc category, such as Pokemon, some of which could appear in the real world in the form of costumes or cuddly toys. The class nature includes 12 outdoor objects, such as rocks or wood. In the category falling, we provide 9 large objects, such as novel trees, that were spawned in an unstable position, which made them fall over. Finally, the airplane class consists of four types of large flying objects.
@@
There are two ways to generate temporal anomalies: Optimization-based or knowledge-based [72]. While both are of value, for our benchmark, it is important to comply with the defined normality, which is more challenging with optimization-based methods. Thus, we provide expert-designed content anomalies. As prior work has shown that expert-defined temporal anomalies require high engineering efforts [11, 12, 2], we provide only a single type of temporal anomalies. We implement sudden braking scenarios of a lead vehicle, which is a very typical scenario in everyday traffic [46]. While we set a planned route for the ego vehicle in our content anomaly scenarios, we use the same route and follow it with a lead vehicle. Then, our ego vehicle follows that lead vehicle. Along the route, the lead vehicle will perform a sudden brake with negative velocities which cannot be seen during training. While braking, we label the lead vehicle in the same way as we label content anomalies.
@@
AnoVox provides 40 labels in the semantic masks, as shown in Fig. 7. We provide labels for standard tasks as used by Cityscapes and CARLA [27, 30]. Additionally, we label the ego vehicle if visible. As described in Sec. 3.2.1, some vehicles in the scene might switch from autopilot to a behavior agent. As this driving behavior is not present in training data, we provide additional labels for all vehicle classes while controlled by a behavior agent. Finally, we provide labels for the included content and temporal anomalies. Labels for the super-classes can be found in the semantic labels, while fine-granular, individual anomaly labels are provided in additional metadata. Finally, voxels are provided with an anomaly label when the closest point to their center is anomalous.
@@ 4.1 Anomaly Detection
To get first insights into how current state-of-the-art methods for content anomalies will perform on our benchmark, we have implemented two candidate methods, one based on camera data and one based on lidar data. For their selection, we have analyzed existing surveys and benchmark results [13, 15, 9, 23]. However, as there is no existing benchmark for lidar-based detection method, this was more challenging. We performed an extensive search, also focusing on open-set segmentation, as this is a similar but more active field. Our final candidates can be found in Table 3. To emphasize the definition of normality based on training data alone, we then neglected all methods that require auxiliary data with anomalies during training. While we are also interested in methods purely trained on raw data, we allowed to usage of semantic labels during training, as the state-of-the-art heavily relies on those. We then picked the best-performing method that provided open-source code for implementations. For camera data, we selected the Rejected By All (RbA) method from Nayal et al. [57], and for lidar data we selected the Redundancy classifier (REAL) framework by Cen et al. [22]. While both methods allow for the usage of auxiliary data during training, we did not fine-tune on anomalies. For RbA this means we did not perform "Outlier Data Exposure" [57], and for REAL this means we only used the "predictive distribution calibration" without "unknown object synthesis" [22].
@@
[…]
Figure 8: Class distributions for the train and evaluation datasets.
@@ 4.1.2 Evaluation
As shown in Table 4, both methods had significant issues detecting anomalies in our setting. While we observed stable training performance and saw improved performance when trained on larger datasets, the results were rather surprising. RbA was able to detect some anomalies successfully, as shown in Fig. 1, but struggled with most. We were able to improve the performance with more training data, but not substantially. REAL on the other hand, while generating well-performing closed-set predictions, was unable to generate any meaningful uncertainties in the absence of unknown objects in the training data, as exemplarily shown in Fig. 9.
This shows that the task of anomaly detection becomes much harder on a challenging benchmark such as AnoVox, where anomalies are not defined by not belonging to Cityscapes classes, but are defined by being absent from the training set, as most state-of-the-art methods currently utilize deviations based on semantic segmentations to derive anomalies. In addition, there is no domain shift for the anomaly classes, as they are rendered in the same simulation engine from which the training data stems from. Finally, lots of frames without visible anomalies and particularly small anomalies raise the bar additionally. This setting, combined with the induced class imbalance due to our quantization loss during voxelization, makes it much harder to perform well on AnoVox compared to existing benchmarks, which we will hope benefit the community greatly in the upcoming years.
@@ 5 Conclusion
In this work, we presented the AnoVox benchmark for ANOmaly detection in autonomous driving. We are the first benchmark to provide ground truth in the form of semantic masks for RGB images and lidar point clouds as well as spatial VOXels, allowing us to compare methods based on arbitrary sensor modalities.
@@
We systematically analyzed the shortcomings of existing benchmarks, most notably no clear definition of normality and a strong focus on RGB data. AnoVox is the first and largest benchmark that provides a concise formalization of normality, accompanied by training data that complies with that definition. Finally, we provided first insights on how well current state-of-the-art models perform on this new paradigm of anomaly detection benchmarking. We see a great need for anomaly detection methods to learn normality based on training data instead of detecting deviations from known Cityscape classes.
Limitations. At the moment, each voxel contains only a single class. This can lead to cases where small anomalies are not represented in single frames of voxelized data. For our evaluation, we map camera-based detection into 3D by using ground truth depth. We have performed experiments using ZoeDepth [7] to map these predictions into 3D, but the overlap between the resulting voxels and the ground truth was too large. Here, more sophisticated methods are necessary. We limited ourselves to only a single temporal anomaly, more types could be implemented. Finally, while simulation is necessary to have full control over both training and evaluation, this leads to a Sim-2-Real gap.
Potential negative impact. As we provide a large number of anomalies, others might use the dataset for extensive fine-tuning. This can lead to cases where anomaly detection methods perform extremely well on benchmarks but might still fail in open-world settings. Thus, we do not recommend using our data with anomalies for anything else than benchmarking.

### v4

@@ Abstract
The scale-up of autonomous vehicles depends heavily on their ability to deal with rare objects on the road. In order to handle such situations, it is necessary to detect anomalies in the first place. Anomaly detection has made great progress in the past years but suffers from poorly designed benchmarks with a strong focus on camera data. In this work, we present AnoVox, the largest benchmark for ANOmaly detection in autonomous driving to date. AnoVox incorporates multimodal sensor data and spatial VOXel ground truth, allowing for the comparison of methods independent of their used sensor. We propose a formal definition of normality and provide a compliant training dataset. AnoVox is the first benchmark to contain both content and temporal anomalies.
@@ 1 Introduction
In the past years, we have seen autonomous vehicles scale from small Operational Design Domains (ODD) to city-wide areas, reaching geographic areas covering multiple neighboring cities in the near future [73, 75, 74]. However, with a growing fleet in an expanding ODD, scenarios from the long tail of rare events [71] occur more frequently as a consequence of increased exposure. Expert perspectives, focusing on content and temporal anomalies, are commonly used to judge "long tail" data points [18, 17, 40, 10, 11, 61, 66], as replaying training data during inference is not possible. Such expert views are often used for the creation of anomaly detection benchmarks but are of no help for anomaly detection. Few publications take a data-oriented perspective on the definition of anomalies [39, 25, 82], also known as out-of-distribution (OOD), corner cases, or outliers. Recent research has focused on anomaly detection methods [42] and its ramifications, such as OOD segmentation [58], object detection [31], instance segmentation [55], and video tracking [54]. These methods typically do not learn a representation of normality from training data alone, but require a semantic segmentation stage and fine-tuning on auxiliary data containing (artificial) anomalies [9, 23]. While there exists a considerable number of datasets in autonomous driving [53, 49, 14], only around 5 - 10% of datasets are designed for anomaly detection [15, 13]. These existing datasets, however, have significant limitations. First, the majority are camera-only, despite autonomous driving relying on multi-modal sensor setups. Second, temporal information is often overlooked, limiting anomaly detection to single frames. Additionally, content anomalies are rarely found in challenging traffic situations, which allows for the assumption that anything on the road might be an anomaly [54]. Most anomalies are human-defined, such as dogs on the street [9], making detection challenging as these classes are typically included in training datasets but not annotated [27, 31]. As a result, anomaly detection methods can miss anomalies they don’t perceive as atypical due to their training.
@@
With AnoVox, we offer a challenging and adaptable benchmark for ANOmaly detection, leveraging the CARLA simulation engine [30] to provide ground truth in all modalities and a spatial VOXel representation. Given the surge in foundation models [21, 60], world models [80, 45], and large-scale pre-training approaches [43, 78], we see potential for the application of self- and unsupervised training paradigms in the field of anomaly detection as well. Hence, we prioritize support for label-free training paradigms and provide an extensive synthetic dataset to enable research in OOD segmentation, addressing the limitations of current datasets. All code is available on GitHub. The dataset can be found on Zenodo. Our contributions include:
@@ 2 Related Work
In autonomous driving, anomaly detection has been classically evaluated on a small set of benchmarks. For our analysis of related works, we have included works with published open-access perception datasets from an ego perspective that provide pixel- or point-wise ground truth. Thus, we neglected frameworks that do not provide explicit data [77, 68, 38, 76, 65], such with missing or incomplete data [34, 52, 48], and any that provide only frame-wise annotations [67, 79, 70, 2]. For our overview of related works as shown in Table 1, we updated and extended the survey of Bogdoll et al. [15].
There are different categories of how anomalies were introduced [15] in the datasets. Recording and Simulation are similar in the way that selected anomalies were directly introduced into the data. This way, the anomalies are truly part of the environment [62, 19, 54]. The definition of what counts as an anomaly can vary, though. Data Augmentation typically follows a copy-and-paste pattern, where images of anomalies are pasted onto a scene from another dataset. This way, a distribution shift between the anomaly and the underlying data is introduced [9, 6]. Web Sourcing describes the process of manually curating images from the web that are deemed anomalous [23, 46]. Class Exclusion is based on existing datasets and removes selected classes from the training data, thus treating them as anomalous, while the classes themselves remain rather normal from a human point of view [35, 41]. Void Classes utilizes void or misc classes from existing datasets and labels them as anomalous. This can be done with additional labeling guidelines [51]. Finally, OOD Detection uses an anomaly detection method to derive anomaly proposals from a dataset which can then be labeled, typically after a human quality inspection [51].
@@
AnoVox is a challenging benchmark that addresses anomaly detection from an embodied AI perspective rather than from a pure computer vision perspective. We provide data in a temporal context with state information about the ego vehicle for typical multimodal sensor setups. For a clear definition of what anomalies are, we provide a formalized definition of normality and compliant training data. This allows for the usage of un- and self-supervised anomaly detection methods, which often rely on large amounts of unlabeled data.
@@ 3.2 Scenario Generation
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 4: Content anomalies: Examples from our six categories: An ape as an animal anomaly, an old tv as a home anomaly, a statue as a special anomaly, a tree as a nature anomaly, a pillar as a falling anomaly, and a hot air balloon as an airplane anomaly.
@@
We provide 178 different content anomalies in five different size classes tiny, small, medium, big, and huge. Semantically, they hierarchically belong to six different super-classes. Every anomaly has an individual label next to its super-class and size label. The class animal includes 33 animals of different sizes. The category home includes 53 typical household items such as furniture, tables, backpacks, or cardboard boxes. The category special includes 67 objects of rather atypical types and such that fall into a misc category, such as Pokemon, some of which could appear in the real world in the form of costumes or cuddly toys. The class nature includes 12 outdoor objects, such as rocks or wood. In the category falling, we provide 9 large objects, such as novel trees, that were spawned in an unstable position, which made them fall over. Finally, the airplane class consists of four types of large flying objects.
@@
There are two ways to generate temporal anomalies: Optimization-based or knowledge-based [72]. While both are of value, for our benchmark, it is important to comply with the defined normality, which is more challenging with optimization-based methods. Thus, we provide knowledge-based content anomalies. As prior work has shown that such expert-defined temporal anomalies require high engineering efforts [11, 12, 2], we provide only a single type of temporal anomalies. We implement sudden braking scenarios of a lead vehicle, which is a very typical scenario in everyday traffic [46]. While we set a planned route for the ego vehicle in our content anomaly scenarios, we use the same route and follow it with a lead vehicle. Then, our ego vehicle follows that lead vehicle. Along the route, the lead vehicle will perform a sudden brake with negative velocities which cannot be seen during training. While braking, we label the lead vehicle in the same way as we label content anomalies.
@@
AnoVox provides 40 label classes, as shown in Fig. 7. We provide labels for standard tasks as used by Cityscapes and CARLA [27, 30]. Additionally, we label the ego vehicle if visible. As described in Sec. 3.2.1, some vehicles in the scene might switch from autopilot to a behavior agent. As this driving behavior is not present in training data, we provide additional labels for all vehicle classes while controlled by a behavior agent. Finally, we provide labels for all content and temporal anomalies. Labels for the super-classes can be found in the semantic labels, while fine-granular, individual anomaly labels are provided in additional metadata. Finally, voxels are provided with an anomaly label when the closest point to their center is anomalous.
@@ 4.1 Anomaly Detection
To get first insights into how current state-of-the-art methods for content anomalies will perform on our benchmark, we have implemented two candidate methods, one based on camera data and one based on lidar data. For their selection, we have analyzed existing surveys and benchmark results [13, 15, 9, 23]. However, as there is no existing benchmark for lidar-based detection method, this was more challenging. We performed an extensive search, also focusing on open-set segmentation, as this is a similar but more active field. Our final candidates can be found in Table 3. To emphasize the definition of normality based on training data alone, we then neglected all methods that require auxiliary data with anomalies during training. While we are also interested in methods purely trained on raw data, we allowed the usage of semantic labels during training, as the state-of-the-art heavily relies on those. We then picked the best-performing method that provided open-source code for implementations. For camera data, we selected the Rejected By All (RbA) method from Nayal et al. [57], and for lidar data we selected the Redundancy classifier (REAL) framework by Cen et al. [22]. While both methods allow for the usage of auxiliary data during training, we did not fine-tune on anomalies. For RbA this means we did not perform "Outlier Data Exposure" [57], and for REAL this means we only used the "predictive distribution calibration" without "unknown object synthesis" [22].
[…]
Figure 8: Class distributions for the train and evaluation datasets.
@@ 4.1.2 Evaluation
As shown in Table 4, both methods had significant issues detecting anomalies. While we observed stable training performance and saw improved performance when trained on larger datasets, the results were rather surprising. RbA was able to detect some anomalies, as shown in Fig. 1, but struggled with most. REAL on the other hand, while generating well-performing closed-set predictions, was unable to generate any meaningful uncertainties in the absence of unknown objects in the training data, as exemplarily shown in Fig. 9.
This shows that the task of anomaly detection becomes much harder on a challenging benchmark such as AnoVox, where anomalies are not defined by not belonging to Cityscapes classes, but are defined by being absent from the training set, as most state-of-the-art methods currently utilize deviations based on semantic segmentations to derive anomalies. In addition, there is no domain shift for the anomaly classes, as they are rendered in the same simulation engine from which the training data stems from. Finally, lots of frames without visible anomalies and particularly small anomalies raise the bar additionally. This setting, combined with the induced class imbalance due to our quantization loss during voxelization, makes it much harder to perform well on AnoVox compared to existing benchmarks, which we hope will benefit the community greatly.
@@ 5 Conclusion
We presented the AnoVox benchmark for ANOmaly detection in autonomous driving. AnoVox provides ground truth in the form of semantic masks for RGB images and lidar point clouds as well as spatial VOXels.
@@
We analyzed the shortcomings of existing benchmarks, most notably no clear definition of normality and a strong focus on RGB data. AnoVox is the first and largest benchmark that provides a concise formalization of normality, accompanied by compliant training data. We provided first insights on how well current SotA models perform on this new paradigm of anomaly detection benchmarking. We see a great need for anomaly detection methods to learn normality based on training data instead of detecting deviations from known Cityscape classes.
Limitations. At the moment, each voxel contains only a single class. This can lead to cases where small anomalies are not represented in single frames of voxelized data. For our evaluation, we map camera-based anomaly detections into 3D by using ground truth depth. We have performed experiments using ZoeDepth [7], a state-of-the-art metric depth estimation model, to map these predictions into 3D, but the overlap between the resulting voxels and the ground truth voxels was too large. Here, more sophisticated methods are necessary. We limited ourselves to only a single temporal anomaly, more types could be implemented. Finally, while simulation is necessary to have full control over both training and evaluation, this leads to a Sim-2-Real gap.
@@
Acknowledgment
This work results from the just better DATA project supported by the German Federal Ministry for Economic Affairs and Climate Action (BMWK), grant number 19A23003H.